OpenStudy (kalli):

can someone help me figure out this question? Find the exact value o the following function. Explain the steps you take to find the value The function is : sec(17pi/6)

OpenStudy (anonymous):

Note, for all trigonometric functions, sec(x) = sex(x + 2pi) Therefore, sec(17pi/6) = sec(5pi/6) sec(5pi/6) = 1 / (cos(5pi/6)) cos(5pi/6) = -rad(3)/2 sec(5pi/6) = -2/rad(3) = -2rad(3) / 3

OpenStudy (anonymous):



OpenStudy (kalli):

thank you @master50777 and thanks for explaining the steps :) I got it now

OpenStudy (anonymous):

Anytime